# Transforming Image Super-Resolution: A ConvFormer-based Efficient Approach

### v1

@@ Abstract
Recent progress in single-image super-resolution (SISR) has achieved remarkable performance, yet the computational costs of these methods remain a challenge for deployment on resource-constrained devices. Especially for transformer-based methods, the self-attention mechanism in such models brings great breakthroughs while incurring substantial computational costs. To tackle this issue, we introduce the Convolutional Transformer layer (ConvFormer) and the ConvFormer-based Super-Resolution network (CFSR), which offer an effective and efficient solution for lightweight image super-resolution tasks. In detail, CFSR leverages the large kernel convolution as the feature mixer to replace the self-attention module, efficiently modeling long-range dependencies and extensive receptive fields with a slight computational cost. Furthermore, we propose an edge-preserving feed-forward network, simplified as EFN, to obtain local feature aggregation and simultaneously preserve more high-frequency information. Extensive experiments demonstrate that CFSR can achieve an advanced trade-off between computational cost and performance when compared to existing lightweight SR methods. Compared to state-of-the-art methods, e.g. ShuffleMixer, the proposed CFSR achieves 0.39 dB gains on Urban100 dataset for ×$\times 2$ SR task while containing 26% and 31% fewer parameters and FLOPs, respectively. Code and pre-trained models are available at https://github.com/Aitical/CFSR.
@@ I Introduction
Single Image Super-Resolution (SISR) is fundamental task in computer vision that aims to enhance the resolution and quality of a single low-resolution image to a higher-resolution image. The goal is to generate a high-resolution image with fine details and improved visual perception from a single input image[1, 2]. The need for SISR arises from various real-world scenarios where high-resolution images are desired but limited by hardware capabilities or constraints. In many applications such as surveillance systems, medical imaging, satellite imagery, and digital photography, the acquisition of high-resolution images may be costly, time-consuming, or restricted. Therefore, SISR techniques provide a valuable solution by leveraging advanced algorithms and computational approaches to upsample low-resolution images.
@@
In recent years, there has been considerable progress in the field of single-image super-resolution (SISR), largely attributed to the advent of deep learning techniques [3, 2]. The primary objective of SISR is to reconstruct a high-resolution image from its low-resolution counterpart, which is a fundamental image processing task in the last decades. A groundbreaking study, SRCNN[4], introduced the concept of learning the mapping between low and high-resolution inputs using a convolutional neural network (CNN). This surpassed the performance of previous methods and led to further exploration of CNN-based methods. Subsequently, several studies have made great development of innovative SR models, such as deep and effective backbones [5, 6, 7, 8] and attention mechanisms [9, 10, 11, 12]. These CNN-based super-resolution methods have significantly advanced the state-of-the-art in SISR and demonstrate the power of deep learning and attention mechanisms in learning complex image representations and generating visually appealing high-resolution images from low-resolution inputs. However, despite the performance gains, these methods typically involve more complex models and higher computational complexity, which hinders their deployment on mobile and edge devices. To address this issue, the design of efficient and lightweight super-resolution models has become crucial. Many works have been proposed to reduce the amounts of parameters or floating point operations (FLOPs) to achieve lightweight models [13, 14, 15, 16, 17, 18, 19, 20, 21]. Zhao et al. [18] proposed the lightweight pixel-attention network (PAN) which replaces the standard residual or dense blocks with an efficient pixel-attention block. Sun et al. [21] proposed ShuffleMixer, which introduces the large kernel convolution into the lightweight SR network and significantly reduces the model complexity by the channel split-shuffle operation.
Recently, Transformer-based architectures have been proposed and attracted great attention in recent years due to their impressive performance[22, 23]. Self-attention mechanism provides the promising long-range modeling and has achieved great breakthroughs in computer vision[22]. While the complexity of it is quadratic to the image size, which requires heavy computational resources. Liu et al.[23] proposed the Swin-Transformer, which performs self-attention within the large local window. Subsequently, many transformer-based SR methods have been advanced [24, 25, 26, 27]. By leveraging self-attention mechanisms and hierarchical architectures, these methods can effectively obtain long-range relations and capture fine details. Transformer-based SR models showcase the promise of performance in addressing image super-resolution challenges that often surpass CNN-based models and achieve new state-of-the-art results. However, the application for lightweight models is limited due to the high computational cost substantial CPU or GPU memory requirements of self-attention mechanism.
@@
In response to the challenge of computational efficiency in image super-resolution, we introduce a novel, self-attention-free approach that offers an excellent balance between computational cost and performance. This positions it as a viable solution for practical applications in lightweight image super-resolution. Specifically, we propose the Convolutional Transformer layer (ConvFormer) as a core component for effective and efficient feature extraction. Building on this foundation, we introduce the ConvFormer-based Super-Resolution network (CFSR) tailored for lightweight image super-resolution tasks. The transformer layer underpinning this approach, as outlined in [22, 23, 28], includes a feature mixer module and a feed-forward network. Drawing inspiration from the recent successes of CNN-based methods [29, 30, 31, 32], our proposed feature mixer module employs large kernel convolutions as gate layers. This innovative design eliminates the need for self-attention in the feature mixer module, efficiently facilitating long-range relations and extensive receptive fields at a minimal additional computational cost. In addition, we introduce an edge-preserving feed-forward network (EFN) that refines the standard feed-forward network by incorporating enhanced edge extraction capabilities. Unlike the conventional feed-forward network (FFN) [22], which incorporates $3\times 3$ depth-wise convolutions for improved local feature aggregation in vision tasks [33, 34], our EFN integrates image gradient priors. This integration not only preserves high-frequency information but also introduces significant improvements for lightweight models without increasing complexity or parameter counts during inference, achieved through re-parameterization [35, 36, 19]. The architecture of CFSR, though straightforward and predominantly convolutional, is significantly more effective than previous methods. When benchmarked against existing methods for the ×$\times 4$ SR task on the Urban100 dataset, as detailed in Figure 1, CFSR demonstrates superior performance. It notably excels in balancing reconstruction quality, model size, and computational efficiency, outperforming state-of-the-art methods with fewer parameters and reduced FLOPs.
@@
    In traditional SISR models, the use of self-attention mechanisms for feature extraction has shown promising results in enhancing super-resolution. However, it comes at a significant computational cost. To address this issue, this paper introduces ConvFormer, a feature mixer based on large kernel convolutions that replaces the self-attention module. ConvFormer efficiently captures long-range dependencies and extensive receptive fields while maintaining a lower computational complexity. This approach demonstrates superior performance and efficiency in lightweight image super-resolution tasks. It will shed new light on the design of CNN-based or hierarchical architecture for lightweight image super-resolution.
@@
    Conventional SISR algorithms often suffer from the loss of high-frequency information, resulting in a degradation of the quality of the generated super-resolved images. To tackle this challenge, this paper proposes EFN, an edge-preserving feed-forward network. EFN incorporates local feature aggregation through convolution layers with edge-preserving filters and preserves high-frequency information using skip connections and deconvolution layers. Compared to traditional methods, EFN achieves better preservation of image details and textures while maintaining high super-resolution performance.
@@
    It presents extensive experiments to verify the effectiveness of CFSR. Compared to the existing advanced methods, CFSR achieves superior performance with less computational cost. Detailed ablation studies are provided to analyze the impact of different components.
In the following section, we will first give some related work of lightweight image super-resolution methods and the progress of modern architectures in Section II. In Section III, we introduce and explain our proposed CFSR method in detail. Then, Section IV describes our training settings and experimental results including ablation analysis, where we compare the performance of our approach to other state-of-the-art methods. Finally, some conclusions are drawn in Section V.
@@ II Related Work
In this section, we will briefly introduce related literature including deep learning-based single-image super-resolution methods, transformer-based architectures, modern convolutional architectures, and the development of re-parameterizing methods.
@@ II-A Single Image Super-Resolution
Recently, deep learning methods have achieved dramatic improvements in SISR tasks [37, 38, 2]. Especially for CNN-based models, various well-designed CNN architectures explore to further improve the SISR performance [5, 39, 6]. VDSR [5] introduces a very deep backbone to predict the residual construction between the LR input and the corresponding HR image. EDSR [6] incorporates residual blocks with skip connections, which allow direct propagation of information from earlier layers to later layers. Besides, attention mechanisms like the channel attention [40] has been introduced to SISR task as well [9, 10, 11, 12]. Zhang et al. [9] introduced the channel attention mechanism and proposed the RCAN model, which extend the backbone into over 400 layers.
In contrast to achieving advancing performance with a rapidly increased number of parameters and computational cost, many lightweight SISR models have been exploited by reducing parameters, especially for resource-limited devices [14, 41, 15, 16, 17, 19, 20]. Hui et al. proposed a deep information distillation network (IDN) [41] and extended it into the information multi-distillation network (IMDN) [15], which won the AIM2020 challenge. Zhang et al. [19] proposed an edge-oriented Convolution Block (ECB) for real-time inference, which extracts 1st-order and 2nd-order spatial derivatives from intermediate features. Most recently, Sun et al. [21] introduced the large kernel convolution into the lightweight SR and proposed the ShuffleMixer. By channel split-shuffle operation, it efficiently reduced latent projection features.
In this paper, we proposed a pretty simple ConvNet for lightweight SR, named CFSR. Compared to the previous work [21], CFSR achieves 0.39 dB gains on Urban100 dataset for ×$\times 2$ SR task while containing 26% and 31% fewer parameter counts and FLOPs, respectively.
@@ II-B Transformer-based Architectures
Most recently, vision transformers have attracted great attention [22, 23] and many work have been proposed to explore transformer-based architectures for image restoration [42, 25, 24, 27, 43]. Various pre-trained models with full or local window-based attention are exploited and applied to target restoration tasks [42, 24, 43, 44]. Liang et al.[25] firstly introduced the Swin-Transformer into the image restoration tasks and proposed a hierarchical architecture SwinIR. Cai et al.[44] proposed the hierarchical patch-based Transformer architecture, which significantly enhances single image super-resolution by progressively recovering high-resolution images through a hierarchy of patch partitions. This multi-stage model begins with small patch sizes, merging them into full resolution in later stages, and includes a unique attention-based position encoding and a multi-receptive field attention module. On the other hand, several works have been studied for lightweight transformer-based models [25, 45, 46, 47]. Wu et al.[45] proposed the lightweight model TCSR which introduces a sliding-window-based self-attention mechanism. One advantage of using transformers for SISR is their ability to capture global context information, which can be beneficial for generating high-quality HR images. However, compared to CNN-based models, transformer-based methods usually requires much more computational resources, even with a small model capacity, such as the SwinIR-light[25].
@@ II-C Modern CNN-based Architectures
Several works investigated some modern CNN-based architectures [32, 31, 48, 29, 49]. On the one hand, large kernel convolution has been revisited[32, 31, 48]. Liu et al.[32] explored a modern CNN-based architecture and introduced larger kernels that utilize $7\times 7$ kernel size. Building upon this work, Ding et al.[31] further brought the kernel size up to 31. Subsequently, Liu et al.[48] further extended the kernel size up to 51 by sparse training. These advancements in kernel size have shown improvements in capturing complex image details and enhancing the super-resolution performance. On the other hand, many work paid more attention to the hierarchical architecture combined convolution and transformer[29, 49]. These architectures leverage the strengths of both convolutional and transformer networks to capture local and global information effectively. Inspired by these findings, in this paper, we exploit a simple transformer-like ConvNet for lightweight SR tasks, where we replace the self-attention module with a large kernel-based mixer and improve the feed-forward network to preserve more high-frequency information. By leveraging the advantages of large kernel convolutions and transformer-like architectures, the proposed method holds potential for achieving better results in SISR while maintaining computational efficiency.
II-D Re-parameterizing Methods
@@
Ding et al.[35] proposed the RepVGG, which provides a practical network architecture and the concept of re-parameterizing operation. By re-parameterizing, multiple linear convolution learned in training process can be merged as one single convolution in inference instead of introducing extra costs. For SISR task, Wang et al.[50] proposed RepSR, a plain architecture for SISR task by re-parameterization. The authors analyzed the impact of the BatchNorm (BN) operation in SISR task and successfully re-introduced BN into SR. Zhang et al.[19] proposed the ECBSR, which introduces more 1st-order and 2nd-order gradient information into the vanilla convolution by re-parameterizing. Wang et al. [51] proposed DDistill-SR, which combines re-parameterization with dynamic convolution[52] to extend the learnable landscape while introducing less complexity in inference.
@@
In this study, we introduce the Convolutional Transformer-based Super-Resolution network (CFSR), a simple yet effective model for lightweight image super-resolution. Drawing inspiration from large kernel methods like ConvNet[32] and re-parameterization strategies exemplified by RepVGG[35], our approach is particularly informed by developments in lightweight super-resolution models such as PAN[18] and ECBSR[19]. CFSR marks a departure from the pixel-attention reliance of PAN, revisiting and streamlining the convolution and self-attention feature extraction mechanisms. We introduce an advanced large kernel feature mixer, engineered to deliver exceptional performance and a significantly expanded receptive field. Additionally, our model capitalizes on the successes of Transformer-based architectures, integrating a novel edge-enhanced re-parameterization operation into the feed-forward network. This innovation not only enhances local feature extraction but also preserves a greater amount of high-frequency information, a significant advancement over the convolution-centric approach of ECBSR[19]. By extending the re-parameterizing branch and focusing on high-frequency detail retention, CFSR stands at the forefront of the current wave of convolutional technique advancements, uniquely tailoring these approaches for the intricate demands of super-resolution tasks.

### v2

@@ Abstract
Recent progress in single-image super-resolution (SISR) has achieved remarkable performance, yet the computational costs of these methods remain a challenge for deployment on resource-constrained devices. In particular, transformer-based methods, which leverage self-attention mechanisms, have led to significant breakthroughs but also introduce substantial computational costs. To tackle this issue, we introduce the Convolutional Transformer layer (ConvFormer) and propose a ConvFormer-based Super-Resolution network (CFSR), offering an effective and efficient solution for lightweight image super-resolution. The proposed method inherits the advantages of both convolution-based and transformer-based approaches. Specifically, CFSR utilizes large kernel convolutions as a feature mixer to replace the self-attention module, efficiently modeling long-range dependencies and extensive receptive fields with minimal computational overhead. Furthermore, we propose an edge-preserving feed-forward network (EFN) designed to achieve local feature aggregation while effectively preserving high-frequency information. Extensive experiments demonstrate that CFSR strikes an optimal balance between computational cost and performance compared to existing lightweight SR methods. When benchmarked against state-of-the-art methods such as ShuffleMixer, the proposed CFSR achieves a gain of 0.39 dB on the Urban100 dataset for the x2 super-resolution task while requiring 26% and 31% fewer parameters and FLOPs, respectively. The code and pre-trained models are available at https://github.com/Aitical/CFSR.
@@ I Introduction
Single Image Super-Resolution (SISR) is a fundamental task in computer vision that aims to enhance the resolution and quality of a low-resolution image, generating a higher-resolution image with finer details and improved visual quality [1, 2]. The need for SISR arises in various real-world scenarios where high-resolution images are desired but are limited by hardware capabilities or constraints. In many applications, such as surveillance systems, medical imaging, satellite imagery, and digital photography—acquiring high-resolution images may be costly, time-consuming, or restricted. Therefore, SISR techniques provide a valuable solution by leveraging advanced algorithms to upsample low-resolution images.
@@
In recent years, there has been considerable progress in SISR, largely attributed to the advent of deep learning techniques [3, 2]. A groundbreaking study, SRCNN [4], introduced the concept of learning the mapping between low- and high-resolution images using convolutional neural networks (CNNs). This surpassed the performance of previous methods and led to further exploration of CNN-based approaches. Subsequently, numerous studies have developed innovative SR models with deep and effective backbones [5, 6] and attention mechanisms [7, 8, 9, 10]. These CNN-based methods have significantly advanced the state-of-the-art in SISR, demonstrating the power of deep learning in learning complex image representations and generating visually appealing high-resolution images from low-resolution inputs. However, despite the performance gains, these methods typically involve more complex models and higher computational complexity, which hinders their deployment on mobile and edge devices. To address this issue, designing efficient and lightweight super-resolution models has become crucial. Many works have been proposed to reduce the number of parameters or floating point operations (FLOPs) to achieve lightweight models [11, 12, 13, 14, 15, 16, 17, 18, 19]. Zhao et al. [16] proposed the lightweight pixel-attention network (PAN), which replaces standard residual or dense blocks with an efficient pixel-attention block. Sun et al. [19] proposed ShuffleMixer, which introduces large kernel convolutions into the lightweight SR network and significantly reduces model complexity through a channel split-shuffle operation. Transformer-based architectures have attracted great attention due to their impressive performance [20, 21]. The self-attention mechanism provides promising long-range modeling capabilities and has achieved significant breakthroughs in computer vision [20]. However, its complexity is quadratic in image size, requiring heavy computational resources. Liu et al. [21] proposed the Swin Transformer, which performs self-attention within large local windows. Subsequently, many transformer-based SR methods have been developed [22, 23, 24, 25, 26, 27]. By leveraging self-attention mechanisms and hierarchical architectures, these methods can effectively capture long-range dependencies and fine details. Transformer-based SR models have often surpassed CNN-based models, achieving new state-of-the-art results in addressing image super-resolution challenges. However, their application in lightweight models is limited due to the high computational cost and substantial CPU or GPU memory requirements of the self-attention mechanism.
@@
To address the challenge of computational efficiency in image super-resolution, we introduce a novel, self-attention-free approach that offers an excellent balance between computational cost and performance, making it a viable solution for practical applications in lightweight image super-resolution. Specifically, we propose the Convolutional Transformer layer (ConvFormer) as a core component for effective and efficient feature extraction. Building on this foundation, we introduce the ConvFormer-based Super-Resolution network (CFSR) tailored for lightweight SISR tasks. As in standard transformer architectures [20, 21, 28], our approach includes a feature mixer module and a feed-forward network. Drawing inspiration from recent successes of CNN-based methods [29, 30, 31, 32], our proposed feature mixer module employs large kernel convolutions as gating layers. This innovative design eliminates the need for self-attention in the feature mixer module, efficiently capturing long-range dependencies and extensive receptive fields with minimal additional computational cost. Additionally, we introduce an edge-preserving feed-forward network (EFN) that refines the standard feed-forward network by incorporating enhanced edge extraction capabilities. Unlike the conventional feed-forward network (FFN) [20], which incorporates $3\times 3$ depth-wise convolutions for improved local feature aggregation in vision tasks [33, 34], our EFN integrates image gradient priors. This integration not only preserves high-frequency information but also introduces significant improvements for lightweight models without increasing complexity or parameter counts during inference, achieved through re-parameterization [35, 36, 17]. The architecture of CFSR, though straightforward and predominantly convolutional, is significantly more effective than previous methods. When benchmarked against existing methods for the ×$\times 4$ SR task on the Urban100 dataset, as detailed in Fig. 1, CFSR demonstrates superior performance. It excels in balancing reconstruction quality, model size, and computational efficiency, outperforming state-of-the-art methods with fewer parameters and reduced FLOPs.
@@
    We introduce ConvFormer, a feature mixer based on large kernel convolutions that replaces the self-attention module in traditional SISR models. ConvFormer efficiently captures long-range dependencies and extensive receptive fields while maintaining lower computational complexity. This approach demonstrates superior performance and efficiency in lightweight image super-resolution tasks, shedding new light on the design of CNN-based or hierarchical architectures for lightweight SISR.
@@
    We propose EFN, an edge-preserving feed-forward network, to address the loss of high-frequency information in conventional SISR algorithms. EFN incorporates local feature aggregation through convolutional layers with edge-preserving filters and preserves high-frequency information using skip connections and deconvolution layers. Compared to traditional methods, EFN achieves better preservation of image details and textures while maintaining high super-resolution performance.
@@
    We present extensive experiments to verify the effectiveness of CFSR. Compared to existing advanced methods, CFSR achieves superior performance with less computational cost. Detailed ablation studies are provided to analyze the impact of different components. Notably, our method outperforms recent NTIRE efficient super-resolution challenge winners.
In the following section, we will first give some related work of lightweight image super-resolution methods and the progress of modern architectures in Sec. II. In Sec. III, we introduce and explain our proposed CFSR method in detail. Then, Sec. IV describes our training settings and experimental results including ablation analysis, where we compare the performance of our approach to other state-of-the-art methods. Finally, some conclusions are drawn in Sec. V.
@@ II Related Work
In this section, we briefly review the related literature, including deep learning-based single-image super-resolution methods, transformer-based architectures, modern convolutional architectures, and the development of re-parameterization methods.
@@ II-A Single Image Super-Resolution
Recently, deep learning methods have achieved dramatic improvements in SISR tasks [37, 38, 2]. In particular, various well-designed CNN architectures have been explored to further improve SISR performance [5, 6, 39, 40, 39, 41]. VDSR [5] introduced a very deep backbone to predict the residual between the low-resolution (LR) input and the corresponding high-resolution (HR) image. EDSR [6] incorporated residual blocks with skip connections, allowing direct propagation of information from earlier layers to later layers. Furthermore, attention mechanisms like channel attention [42] have been introduced to the SISR task [7, 8, 9, 10]. Zhang et al. [7] proposed the RCAN model, which extends the backbone to over 400 layers by incorporating channel attention mechanisms.
In contrast to achieving advanced performance with a rapidly increased number of parameters and computational cost, many lightweight SISR models have been developed to reduce parameters, especially for resource-limited devices [11, 12, 43, 13, 14, 15, 17, 18]. Hui et al. proposed the Deep Information Distillation Network (IDN) [43] and extended it to the Information Multi-Distillation Network (IMDN) [13], which won the AIM2020 challenge. Zhang et al. [17] proposed the Edge-oriented Convolution Block (ECB) for real-time inference, which extracts first-order and second-order spatial derivatives from intermediate features.
Recent NTIRE challenges have driven significant advancements in efficient super-resolution. Liu et al. introduced the Residual Local Feature Network (RLFN) [44] in NTIRE 2022, employing large kernel convolutions and deformable convolutions to efficiently capture long-range dependencies. This approach demonstrated the potential of adapting techniques from medical image segmentation to the SR task. In NTIRE 2023, Xu et al. proposed the Deep Image Prior Network (DIPNet) [45], introducing a multi-stage lightweight network boosting method. DIPNet leverages enhanced high-resolution output as additional supervision and employs network simplification techniques like re-parameterization and iterative pruning.
Most recently, Sun et al. [19] introduced large kernel convolutions into lightweight SR and proposed ShuffleMixer. Through a channel split-shuffle operation, it efficiently reduces latent projection features. Building on this trend, Zhang et al. developed the Swift Parameter-free Attention Network (SPAN) [46] for NTIRE 2024, which introduces a novel parameter-free attention mechanism. SPAN uses symmetric activation functions and residual connections to enhance high-contribution information while suppressing redundant information, achieving a significant quality-speed trade-off.
@@ II-B Transformer-based Architectures
Recently, vision transformers have attracted great attention [20, 21], and many works have been proposed to explore transformer-based architectures for image restoration [47, 23, 24, 22, 27, 25]. Various pre-trained models with full or local window-based attention have been exploited and applied to target restoration tasks [47, 22, 25, 48]. Liang et al. [23] first introduced the Swin Transformer into image restoration tasks and proposed the hierarchical architecture SwinIR. Cai et al. [48] proposed the Hierarchical Patch-based Transformer architecture, which significantly enhances single-image super-resolution by progressively recovering high-resolution images through a hierarchy of patch partitions. Recognizing the importance of locality in vision tasks, Li et al. proposed LocalViT [49], which introduces depth-wise convolution into the feed-forward network of vision transformers. This approach aims to combine the global modeling capabilities of transformers with the local feature extraction strengths of CNNs. LocalViT demonstrated significant performance improvements over baseline models with minimal increases in parameters and computational cost. Several works have focused on lightweight transformer-based models [23, 50, 51, 52]. Wu et al. [50] proposed the lightweight model TCSR, which introduces a sliding-window-based self-attention mechanism. One advantage of using transformers for SISR is their ability to capture global context information, beneficial for generating high-quality HR images. However, compared to CNN-based models, transformer-based methods usually require much more computational resources, even with a small model capacity, such as SwinIR-light [23].
@@ II-C Modern CNN-based Architectures
Several works have investigated modern CNN-based architectures [32, 31, 53, 29, 54]. On one hand, large kernel convolutions have been revisited [32, 31, 53]. Liu et al. [32] explored a modern CNN-based architecture and introduced larger kernels utilizing a $7\times 7$ kernel size. Building upon this work, Ding et al. [31] increased the kernel size up to 31. Subsequently, Liu et al. [53] extended the kernel size up to 51 through sparse training. These advancements in kernel size have shown improvements in capturing complex image details and enhancing super-resolution performance.
@@
On the other hand, many works have focused on hierarchical architectures combining convolutional and transformer elements [29, 54]. These architectures leverage the strengths of both convolutional and transformer networks to effectively capture local and global information. Inspired by these findings, we exploit a simple transformer-like ConvNet for lightweight SR tasks, where we replace the self-attention module with a large kernel-based mixer and improve the feed-forward network to preserve more high-frequency information. By leveraging the advantages of large kernel convolutions and transformer-like architectures, the proposed method holds potential for achieving better results in SISR while maintaining computational efficiency.
II-D Re-parameterization Methods
@@
Ding et al. [35] proposed RepVGG, which provides a practical network architecture and the concept of re-parameterization. By re-parameterizing, multiple linear convolutions learned during training can be merged into a single convolution during inference without introducing extra computational costs. For the SISR task, Wang et al. [55] proposed RepSR, a plain architecture for SISR via re-parameterization. The authors analyzed the impact of the Batch Normalization (BN) operation in SISR and successfully reintroduced BN into SR. Zhang et al. [17] proposed ECBSR, which introduces more first-order and second-order gradient information into the vanilla convolution via re-parameterization. Wang et al. [56] proposed DDistill-SR, which combines re-parameterization with dynamic convolution [57] to extend the learnable landscape while introducing less complexity in inference.
@@
In this study, we introduce the Convolutional Transformer-based Super-Resolution network (CFSR), a simple yet effective model for lightweight image super-resolution. Drawing inspiration from large kernel methods like ConvNet [32] and re-parameterization strategies exemplified by RepVGG [35], our approach is particularly informed by developments in lightweight super-resolution models such as PAN [16] and ECBSR [17]. CFSR marks a departure from the pixel-attention reliance of PAN, revisiting and streamlining convolution and self-attention feature extraction mechanisms. We introduce an advanced large kernel feature mixer, engineered to deliver exceptional performance and a significantly expanded receptive field. Additionally, our model capitalizes on the successes of transformer-based architectures, integrating a novel edge-enhanced re-parameterization operation into the feed-forward network. This innovation not only enhances local feature extraction but also preserves a greater amount of high-frequency information, a significant advancement over the convolution-centric approach of ECBSR [17]. By extending the re-parameterization branch and focusing on high-frequency detail retention, CFSR stands at the forefront of the current wave of convolutional technique advancements, uniquely tailoring these approaches for the intricate demands of super-resolution tasks.